A rectangular vegetable garden has a perimeter of 64ft. the length ot the garden is 20ft. what is the width of the garden.
P= 2l + 2w
64ft=2(20) + 2w
64ft= 40 + 2w
-40 -40
24= 2w
24/2= 2w/2
12= w

Answer by gen2011(1)

A rectangular vegetable garden has a perimeter of 64ft. the length of the garden is 20ft. what is the width of the garden.

P= 2l + 2w
64ft=2(20) + 2w
64ft= 40 + 2w
-40 -40
24= 2w
24/2= 2w/2
12= w